# ] 1NLPR, MAIS, CASIA  2UCAS  3ByteDance

### v1

@@ 1 Introduction
Figure 1: Qualitative examples from TreeBench for each discipline. Each question requires focused visual parsing on mere objects, and some even request second-order reasoning beyond precise localization. Moreover, the bounding boxes of all target instances are provided, ensuring a traceable evaluation. All these questions are challenging, as OpenAI-o3 [43] and Gemini-2.5-Pro [10] cannot answer them correctly simultaneously.
Recent breakthroughs in Large Language Models (LLMs) reasoning, such as OpenAI-o1 [42] and DeepSeek-R1 [15] with remarkable test-time scaling properties, have motivated researchers to explore reasoning for Large Multimodal Models (LMMs) [18, 64, 65, 6]. These models are typically remarkable in their mathematical and scientific reasoning, particularly through text-space reasoning. However, they exhibit critical limitations when applied to perception-heavy tasks [19] or general multimodal benchmarks [62], primarily due to accumulated language bias from their exclusive reliance on textual reasoning pathways. A paradigm shift toward visual grounded reasoning emerged with models like OpenAI-o3 [43], which is able to “think with images” by dynamically referencing and amplifying task-relevant regions during reasoning, resulting in image-text interleaved reasoning pathways. Yet, despite growing interest, the community currently lacks comprehensive evaluation benchmarks for assessing these capabilities.
Classical benchmarks like POPE [26], MMBench [30], SEED-Bench [24], and MMMU [71] usually overlook fine-grained localization and verifiable reasoning chains. Others [66, 72, 63, 11, 56, 55, 73] partially address localization but lack traceability or complex reasoning: V* Bench [66] is restricted to simple spatial queries (e.g., “Is A left of B?”) and risks data contamination with COCO-derived images [27]; MME-RealWorld [72], HR-Bench [63], and document benchmarks [4, 39, 31] support high-resolution inputs but lack traceable evidence and second-order reasoning such as perspective shifts. In short, these benchmarks fail to adequately evaluate three key elements central to visual grounded reasoning: nuanced visual grounding, traceable multi-step reasoning, and dynamic cross-modal interaction through interleaved box-text reasoning pathways.
@@
To conduct TreeBench, we sample 1K images from SA-1B [20], prioritizing images with dense objects, as SA-1B [20] offers high-resolution, real-world scenes with a large number of small and varied objects, making it particularly suitable for evaluating visual grounded reasoning. Subsequently, 8 experts with solid technical backgrounds are involved in hand-crafted annotation for 10 sub-tasks, as demonstrated in Figure 1. In particular, we present a semi-automated pipeline. Each of OpenAI-o3 [43] and Gemini-2.5-Pro [10] is required to create three distinct questions belonging to a specific subtask, accompanied by multiple-choice options and the respective correct answers. Subsequently, experts curated or replaced these to ensure quality and difficulty. We additionally incorporate a cross-verification stage for further quality control. Finally, TreeBench incorporates 405 high-quality and extremely challenging VQA pairs with accurate bounding boxes of target instances. A comprehensive comparison between our TreeBench and other related benchmarks is provided in Table 1. Key advantages are summarized as follows:
  • Annotation Quality. Unlike benchmarks relying on LMM-generated labels such as MMT-Bench [69] and SEED-Bench [24], our expert-driven process ensures correctness and extreme difficulty. However, relying on models would inevitably introduce significant noise, compromising the quality of the annotations. On the contrary, our TreeBench is manually designed by 8 LMM experts, ensuring the annotation correctness and ensuring the difficulty of all questions.
@@
  • Task Difficulty. While models approach saturation (>90%) on benchmarks like V* Bench [66], even open-sourced state-of-the-art performers like Qwen2.5-VL-72B [2] achieve only 42.2 on our TreeBench, implying a large potential improvement for future works.
@@
V* Bench [66] 2,246×1,583 85.9
HR-Bench-4K [63] 4,023×3,503 79.3
HR-Bench-8K [63] 5,727×4,430 76.0
MME-RealWorld [72] 2,076×1,434 62.9
@@
Table 1: Comparison between benchmarks related to “thinking with images”. TreeBench features traceable evidence annotations, as well as high input resolution and challenging questions. $\downarrow$ indicates the lower, the more challenging.
Figure 2: Normalized performance comparison with our TreeVGR and other works [2, 76, 50] on our TreeBench for each category.
Beyond evaluation, we further introduce TreeVGR (Traceable Evidence for Visual Grounded Reasoning), a training paradigm enhancing localization-driven visual reasoning. Previous attempts like [60, 76, 5, 13, 46, 44, 50, 32] solely supervise final answers and neglect intermediate region-of-interest generation processes. It becomes hard to quantify the actual contribution of the “grounding-then-answering” framework. On the contrary, we propose TreeVGR, a novel training methodology emphasizing traceable evidence through reinforcement learning (RL), which explicitly supervises bounding box generation.
Building on RL with conventional accuracy-based and formatting rewards, TreeVGR leverages a novel dual IoU reward to ensure both precision and recall in localizing the ground-truth bounding boxes for each target instance. To implement this, we curate 37K samples for RL training, each comprising an image, a question, an answer, and corresponding bounding box annotations for all target instances. Empirically, initialized from Qwen2.5-VL-7B [2], TreeVGR brings significant improvements on various benchmarks, i.e., +16.8 on V* Bench [66], +12.6 on MME-RealWorld-Lite [72], and +13.4 on our TreeBench. Moreover, as illustrated in Figure 2, compared with related approaches, our TreeVGR enables traceable and explainable reasoning pathways with more accurate localizations (mIoU), and finally contributes to bootstrapped overall performance.
@@ 2 Related Work
Large Multimodal Models. Initial breakthroughs in Large Multimodal Models (LMMs), such as Flamingo [1] and BLIP-2 [25], achieved this by integrating visual features into the LLM backbone via cross-attention mechanisms. A significant shift towards efficiency emerged with LLaVA [28], which introduced a much more efficient approach. It projects visual features from a pre-trained encoder (e.g., CLIP [45]) directly into the LLM’s semantic space using a simple two-layer MLP. This paradigm of feature projection catalyzed rapid advancement. Subsequent research has dramatically scaled LMM capabilities and tackled increasingly complex tasks [57, 29, 23, 58, 2, 61, 22, 77, 67, 59]. A critical frontier has been handling high-resolution inputs. Models like LLaVA-NeXT [29] and InternVL-1.5 [8] adopt any resolution strategy. Qwen2-VL [61] and Qwen2.5-VL [2] introduce multimodal Rotary Position Embedding (mROPE) to support arbitrary resolution inputs. Beyond resolution, scaling pretraining with high-quality data is also vital, as demonstrated by InternVL3 [77]. Collectively, these models represent the state-of-the-art, forming robust baselines for diverse real-world multimodal applications. Our work builds upon these advances by leveraging their strong native visual grounding capabilities. However, existing LMMs do not naturally perform an explicit "grounding-then-answering" process, often resulting in misaligned or incomplete responses. By explicitly modeling this sequential process, our approach ensures more accurate and interpretable answers through grounded reasoning.
@@
Reasoning LMMs. The groundbreaking reasoning capabilities of LLMs, exemplified by systems like OpenAI-o1 [42] and DeepSeek-R1 [15] have motivated efforts to extend similar competencies to multimodal settings using reinforcement learning (RL) [51]. Early approaches primarily focused on equipping LMMs to solve complex math and science problems involving image inputs [18, 64, 65, 6]. Other approaches [48, 33, 3] directly adopt GRPO [47] to open-ended visual grounding. Moreover, some attempts [34, 40, 46, 44] focus on regions-of-interest localization before actually answering the question. A recent milestone, OpenAI-o3 [43], advanced multimodal reasoning by enabling dynamic image manipulation, e.g., cropping and zooming into regions of interest, to emulate human-like "thinking with images." Subsequent research has sought to replicate this capability through diverse strategies: constructing SFT data [60], vanilla RL [13], framing grounding as a function [76], decoupling grounding and answering [5], multi-task reinforcement learning [32], and curiosity-driven reasoning [50]. Critically, these RL-based methods supervise only the final answer. In contrast, our TreeVGR emphasizes traceable evidence during RL training, i.e., supervising generated bounding boxes to ensure precise localization throughout the reasoning process. By doing so, TreeVGR enables more transparent, reliable, and fine-grained control over the reasoning pipeline.
@@
Benchmarks for LMMs. Current benchmarks lack comprehensive evaluation of multimodal models’ ability to “think with images”, a capability demanding three core competencies: (1) focused visual perception (identifying small targets in large scenes), (2) traceable evidence (evaluating generated bounding boxes for explainability), and (3) second-order reasoning (deriving insights beyond precise instance localization). Some benchmarks may partially satisfy the first condition. While some benchmarks address isolated aspects, critical gaps persist. Classical benchmarks like POPE [26], MMBench [30], SEED-Bench [24], and MMMU [71] usually overlook fine-grained localization and verifiable reasoning chains. V* [66] evaluates detailed attributes and spatial relationships (e.g., “Is A left of B?”) but relies on COCO-derived images [27], introducing high contamination risk. MME-RealWorld [72] and HR-Bench [63] support high-resolution inputs but lack traceable evidence, and their questions often become easy when grounded precisely. Crucially, no benchmark integrates all three requirements, particularly the need for complex reasoning conditional on precise grounding, e.g., perspective transform: “From the perspective of person A, what is the relative direction of object B?”. To bridge this gap, we propose TreeBench, the first benchmark designed explicitly for “thinking with images” with traceable, multistep evaluation. Beyond accuracy, TreeBench assesses: (1) region quality, i.e., faithfulness of generated regions-of-interest in visual reasoning chains, and (2) second-order reasoning, i.e., capabilities requiring inference beyond localization. State-of-the-art models, Gemini-2.5-Pro [10] and OpenAI-o3 [43], perform poorly on TreeBench (<60%), underscoring its rigor and the unmet challenges in multimodal reasoning.
@@ 3 TreeBench
TreeBench is designed to address a critical gap in multimodal evaluation by establishing the first comprehensive benchmark for assessing “thinking with images” capabilities. Specifically, it mainly evaluates (1) the ability of identifying small target objects with long, detailed, and unique text captions in large, complex, and real-world scenes, (2) the explainability of reasoning pathways and traceable evidence, and (3) second-order reasoning beyond precise localization. Our TreeBench systematically evaluates 10 core competencies through 405 distinct questions, organized into two progressive protocols, i.e., “Perception” and “Reasoning”, with representative examples demonstrated in Figure 1.
3.1 Task Definition
Perception evaluates the model’s ability to accurately “see” and “identify” specific content, which is one of the basic capabilities of directly extracting and interpreting visual information from every detail of the provided image. These tasks primarily evaluate first-order visual reasoning capabilities, where correct answers usually depend on the accurate localization of target questions (e.g., objects, regions, or text) and directly recognize their explicit attributes without requiring higher-level logical inference or abstract conceptualization. The success hinges on precise perceptual grounding. It includes:
@@ 3.1 Task Definition
3.2 Annotation Pipeline
The TreeBench dataset was constructed through a systematic pipeline combining automated sampling, LMM-assisted generation, and three rounds of human validation. The annotation team contains eight human experts in LMMs, including six Ph.D candidates and two senior research scientists.
Image Selection. A total of 1K images are initially sampled from the SA-1B [20], with deliberate prioritization of images containing high-density objects (e.g., scenes with overlapping or clustered items), as SA-1B [20] offers high-resolution, real-world scenes with a large number of small and varied objects, making it particularly suitable for evaluating visual grounded reasoning. To ensure balanced representation across categories, 100 images are initially allocated per category.
First Round Quality Control. The annotation team manually evaluates the relevance and quality of each image for its assigned category. This step is critical for addressing category-specific requirements, e.g., the “Ordering” category necessitates images with visually similar or repetitive objects for practical reasoning tasks. Following this review, 647 images meet the criteria for inclusion in subsequent stages.
Automated Question Generation. Question-option-answer trios are then generated using two advanced LMMs, i.e., OpenAI-o3 [43] and Gemini-2.5-Pro [10], each tasked with producing three diverse, high-quality questions per image. Prompts are designed to emphasize task-specific complexity and visual-semantic alignment.
Second Round Quality Control. Human experts then manually review all six model-generated questions per image. For each image, annotators selected the most semantically coherent and task-relevant question from the pool of six, prioritizing: (1) alignment with the target subtask, (2) avoidance of trivial or ambiguous object referring, and (3) clarity and unambiguous answerability. If none of the six questions met these criteria, annotators manually constructed a new question tailored to the image. This step ensured that only high-quality, human-vetted questions advanced to the next stage.
Difficulty Filtering. Questions deemed insufficiently challenging are removed through model-based consensus screening. Specifically, any question answered correctly by all four state-of-the-art vision-language models (Qwen2.5-VL-72B [2], InternVL3-78B [77], GPT-4o [41], Gemini-2.5-Flash [9]) was excluded to ensure the benchmark retained meaningful difficulty.
Third Round Quality Control. The final cross-verification phase engages independent human annotators to cross-validate the accuracy and relevance of each question-option-answer pair. The final dataset comprised 405 rigorously validated questions.
3.3 Statistics
Distribution of Target Instance Area. We compute the relative area for each target instance using its bounding box, i.e., $\text{area}=\frac{1}{HW}(y_{2}-y_{1})(x_{2}-x_{1})$, where $H$ and $W$ are the input resolution. Figure 6 is the histogram of the mean area for each question. It illustrates that the majority of target instances in TreeBench are extremely small, with a sharp peak near 0.0 and a long tail extending to larger areas (up to 0.7). The mean area across all questions is 0.0305, confirming that targets are predominantly tiny. Most questions (highest frequency bin) involve target instances with areas clustered around 0.0 to 0.05, while only a small fraction require identifying larger objects. This distribution highlights the importance of addressing challenging scenarios where small-scale object detection and reasoning are crucial, potentially compromising model performance due to limited visual information.

### v2

# Traceable Evidence Enhanced Visual Grounded Reasoning: Evaluation and Method

@@ 1 Introduction
Figure 1: Qualitative examples from TreeBench for each discipline. Each question requires focused visual parsing on mere objects, and some even request second-order reasoning beyond precise localization. Moreover, the bounding boxes of all target instances are provided, ensuring a traceable evaluation. All these questions are challenging, as OpenAI-o3 (OpenAI, 2025) and Gemini-2.5-Pro (DeepMind, 2025b) cannot answer them correctly simultaneously.
Recent breakthroughs in Large Language Models (LLMs) reasoning, such as OpenAI-o1 (OpenAI, 2024b) and DeepSeek-R1 (Guo et al., 2025a) with remarkable test-time scaling properties, have motivated researchers to explore reasoning for Large Multimodal Models (LMMs) (Huang et al., 2025; Wei et al., 2025a; b; Chen et al., 2025). These models are typically remarkable in their mathematical and scientific reasoning, particularly through text-space reasoning. However, they exhibit critical limitations when applied to perception-heavy tasks (Jiang et al., 2025) or general multimodal benchmarks (Wang et al., 2024c), primarily due to accumulated language bias from their exclusive reliance on textual reasoning pathways. A paradigm shift toward visual grounded reasoning emerged with models like OpenAI-o3 (OpenAI, 2025), which is able to “think with images” by dynamically referencing and amplifying task-relevant regions during reasoning, resulting in image-text interleaved reasoning pathways. Yet, despite growing interest, the community currently lacks comprehensive evaluation benchmarks for assessing these capabilities.
Classical benchmarks like POPE (Li et al., 2023c), MMBench (Liu et al., 2023b), SEED-Bench (Li et al., 2023a), and MMMU (Yue et al., 2024) usually overlook fine-grained localization and verifiable reasoning chains. Others (Wu and Xie, 2024; Zhang et al., 2024a; Wang et al., 2025f; Dong et al., 2024; Wang et al., 2025b; a; Zhang et al., 2024b) partially address localization but lack traceability or complex reasoning: V* Bench (Wu and Xie, 2024) is restricted to simple spatial queries (e.g., “Is A left of B?”) and risks data contamination with COCO-derived images (Lin et al., 2014); MME-RealWorld (Zhang et al., 2024a), HR-Bench (Wang et al., 2025f), and document benchmarks (Biten et al., 2022; Mathew et al., 2021; Liu et al., 2023c) support high-resolution inputs but lack traceable evidence and second-order reasoning such as perspective shifts. In short, these benchmarks fail to adequately evaluate three key elements central to visual grounded reasoning: nuanced visual grounding, traceable multi-step reasoning, and dynamic cross-modal interaction through interleaved box-text reasoning pathways.
@@
To conduct TreeBench, we sample 1K images from SA-1B (Kirillov et al., 2023), prioritizing images with dense objects, as SA-1B (Kirillov et al., 2023) offers high-resolution, real-world scenes with many small and varied objects, making it particularly suitable for evaluating visual grounded reasoning. Subsequently, 8 experts with solid technical backgrounds are involved in hand-crafted annotation for 10 sub-tasks, as demonstrated in Figure 1. In particular, we present a semi-automated pipeline. Each of OpenAI-o3 (OpenAI, 2025) and Gemini-2.5-Pro (DeepMind, 2025b) is required to create three distinct questions belonging to a specific subtask, accompanied by multiple-choice options and the respective correct answers. Subsequently, experts curated or replaced these to ensure quality and difficulty. We additionally incorporate a cross-verification stage for further quality control. Finally, TreeBench incorporates 405 high-quality and extremely challenging VQA pairs with accurate bounding boxes of target instances. A comprehensive comparison between TreeBench and other related benchmarks is provided in Table 1. Key advantages are:
  • Annotation Quality. Unlike benchmarks relying on LMM-generated labels such as MMT-Bench (Ying et al., 2024) and SEED-Bench (Li et al., 2023a), our expert-driven process ensures correctness and extreme difficulty. However, relying on models would inevitably introduce significant noise, compromising the quality of the annotations. On the contrary, our TreeBench is manually designed by 8 LMM experts, ensuring the annotation correctness and ensuring the difficulty of all questions.
@@
  • Task Difficulty. While models approach saturation (>90%) on benchmarks like V* Bench (Wu and Xie, 2024), even open-sourced state-of-the-art performers like Qwen2.5-VL-72B (Bai et al., 2025a) achieve only 42.2 on our TreeBench, implying a large potential improvement.
Table 1: Comparison between benchmarks related to “thinking with images”. TreeBench features traceable evidence annotations, as well as high input resolution and challenging questions.
@@
V* Bench 2,246×1,583 85.9
HR-Bench-4K 4,023×3,503 79.3
HR-Bench-8K 5,727×4,430 76.0
MME-RealWorld 2,076×1,434 62.9
@@
Figure 2: Normalized performance comparison with our TreeVGR and other works (Bai et al., 2025a; Zheng et al., 2025b; Su et al., 2025) on our TreeBench for each category.
Beyond evaluation, we further introduce TreeVGR (Traceable Evidence for Visual Grounded Reasoning), a training paradigm enhancing localization-driven visual reasoning. Previous attempts like (Wang et al., 2025e; Zheng et al., 2025b; Cao et al., 2025; Fan et al., 2025; Shao et al., 2024a; Qi et al., 2024; Su et al., 2025; Liu et al., 2025a) solely supervise final answers and neglect intermediate region-of-interest generation processes. It becomes hard to quantify the actual contribution of the “grounding-then-answering” framework. On the contrary, we propose TreeVGR, a novel training methodology emphasizing traceable evidence through reinforcement learning (RL), which explicitly supervises bounding box generation.
Building on RL with conventional accuracy-based and formatting rewards, TreeVGR leverages a novel dual IoU reward to ensure both precision and recall in localizing the ground-truth bounding boxes for each target instance. To implement this, we curate 37K samples for RL training, each comprising an image, a question, an answer, and corresponding bounding box annotations for all target instances. Empirically, initialized from Qwen2.5-VL-7B (Bai et al., 2025a), TreeVGR brings significant improvements on various benchmarks, i.e., +16.8 on V* Bench (Wu and Xie, 2024), +12.6 on MME-RealWorld-Lite (Zhang et al., 2024a), and +13.4 on our TreeBench. Moreover, as illustrated in Figure 2, compared with related approaches, our TreeVGR enables traceable and explainable reasoning pathways with more accurate localizations (mIoU), and finally contributes to bootstrapped overall performance.
@@ 2 Related Work
Large Multimodal Models. Initial breakthroughs in Large Multimodal Models (LMMs), such as Flamingo (Alayrac et al., 2022) and BLIP-2 (Li et al., 2023b), achieved this by integrating visual features into the LLM backbone via cross-attention mechanisms. A significant shift towards efficiency emerged with LLaVA (Liu et al., 2023a), which introduced a much more efficient approach. It projects visual features from a pre-trained encoder (e.g., CLIP (Radford et al., 2021)) directly into the LLM’s semantic space using a simple two-layer MLP. This paradigm of feature projection catalyzed rapid advancement. Subsequent research has dramatically scaled LMM capabilities and tackled increasingly complex tasks (Wang et al., 2025c; Liu et al., 2024a; Li et al., 2024; Wang et al., 2025d; Bai et al., 2025a; Wang et al., 2024b; Lei et al., 2025; Zhu et al., 2025; Wu et al., 2024; Wang et al., 2024a; Yang et al., 2024a; b; 2025b; 2025a; 2026). A critical frontier has been handling high-resolution inputs. Models like LLaVA-NeXT (Liu et al., 2024a) and InternVL-1.5 (Chen et al., 2024b) adopt any resolution strategy. Qwen2-VL (Wang et al., 2024b) and Qwen2.5-VL (Bai et al., 2025a) introduce multimodal Rotary Position Embedding (mROPE) to support arbitrary resolution inputs. Beyond resolution, scaling pretraining with high-quality data is also vital, as demonstrated by InternVL3 (Zhu et al., 2025). Collectively, these models represent the state-of-the-art, forming robust baselines for diverse real-world multimodal applications. Our work builds upon these advances by leveraging their strong native visual grounding capabilities. However, existing LMMs do not naturally perform an explicit "grounding-then-answering" process, often resulting in misaligned or incomplete responses. By explicitly modeling this sequential process, our approach ensures more accurate and interpretable answers through grounded reasoning.
@@
Reasoning LMMs. The groundbreaking reasoning capabilities of LLMs, exemplified by systems like OpenAI-o1 (OpenAI, 2024b) and DeepSeek-R1 (Guo et al., 2025a) have motivated efforts to extend similar competencies to multimodal settings using reinforcement learning (RL) (Sutton et al., 1998). Early approaches primarily focused on equipping LMMs to solve complex math and science problems involving image inputs (Huang et al., 2025; Wei et al., 2025a; b; Chen et al., 2025). Other approaches (Shen et al., 2025; Liu et al., 2025b; Bai et al., 2025b) directly adopt GRPO (Shao et al., 2024b) to open-ended visual grounding. Moreover, some attempts (Liu et al., 2024b; Mondal et al., 2024; Shao et al., 2024a; Qi et al., 2024) focus on regions-of-interest localization before actually answering the question. A recent milestone, OpenAI-o3 (OpenAI, 2025), advanced multimodal reasoning by enabling dynamic image manipulation, e.g., cropping and zooming into regions of interest, to emulate human-like "thinking with images." Subsequent research has sought to replicate this capability through diverse strategies: constructing SFT data (Wang et al., 2025e), vanilla RL (Fan et al., 2025), framing grounding as a function (Zheng et al., 2025b), decoupling grounding and answering (Cao et al., 2025), multi-task reinforcement learning (Liu et al., 2025a), and curiosity-driven reasoning (Su et al., 2025). Critically, these RL-based methods supervise only the final answer. In contrast, our TreeVGR emphasizes traceable evidence during RL training, i.e., supervising generated bounding boxes to ensure precise localization throughout the reasoning process. By doing so, TreeVGR enables more transparent, reliable, and fine-grained control over the reasoning pipeline.
@@
Benchmarks for LMMs. Current benchmarks lack comprehensive evaluation of multimodal models’ ability to “think with images”, a capability demanding three core competencies: (1) focused visual perception (identifying small targets in large scenes), (2) traceable evidence (evaluating generated bounding boxes for explainability), and (3) second-order reasoning (deriving insights beyond precise instance localization). Some benchmarks may partially satisfy the first condition. While some benchmarks address isolated aspects, critical gaps persist. Classical benchmarks like POPE (Li et al., 2023c), MMBench (Liu et al., 2023b), SEED-Bench (Li et al., 2023a), and MMMU (Yue et al., 2024) usually overlook fine-grained localization and verifiable reasoning chains. V* (Wu and Xie, 2024) evaluates detailed attributes and spatial relationships (e.g., “Is A left of B?”) but relies on COCO-derived images (Lin et al., 2014), introducing high contamination risk. MME-RealWorld (Zhang et al., 2024a) and HR-Bench (Wang et al., 2025f) support high-resolution inputs but lack traceable evidence, and their questions often become easy when grounded precisely. Crucially, no benchmark integrates all three requirements, particularly the need for complex reasoning conditional on precise grounding, e.g., perspective transform: “From the perspective of person A, what is the relative direction of object B?”. To bridge this gap, we propose TreeBench, the first benchmark designed explicitly for “thinking with images” with traceable, multistep evaluation. Beyond accuracy, TreeBench assesses: (1) region quality, i.e., faithfulness of generated regions-of-interest in visual reasoning chains, and (2) second-order reasoning, i.e., capabilities requiring inference beyond localization. State-of-the-art models, Gemini-2.5-Pro (DeepMind, 2025b) and OpenAI-o3 (OpenAI, 2025), perform poorly on TreeBench (<60%), underscoring its rigor and the unmet challenges in multimodal reasoning.
@@ 3 TreeBench
TreeBench is designed to address a critical gap in multimodal evaluation by establishing the first comprehensive benchmark for assessing “thinking with images” capabilities. Specifically, it mainly evaluates (1) the ability of identifying small target objects with long, detailed, and unique text captions in large, complex, and real-world scenes, (2) the explainability of reasoning pathways and traceable evidence, and (3) second-order reasoning beyond precise localization. Our TreeBench systematically evaluates 10 core competencies through 405 distinct questions, organized into two progressive protocols, i.e., “Perception” and “Reasoning”, with representative examples in Figure 1. In the following, we provide a detailed exploration of task definitions. The annotation pipeline and the final statistics of TreeBench can be found in Appendix B and Appendix C, respectively.
Perception evaluates the model’s ability to accurately “see” and “identify” specific content, which is one of the basic capabilities of directly extracting and interpreting visual information from every detail of the provided image. These tasks primarily evaluate first-order visual reasoning capabilities, where correct answers usually depend on the accurate localization of target questions (e.g., objects, regions, or text) and directly recognize their explicit attributes without requiring higher-level logical inference or abstract conceptualization. It includes: